OpenStudy (anonymous):

Add and answer in simplest fractional form. 3/8 + -1/8 A. 1/8 B. 2/16 C. 1/4

sam (.sam.):

1/4 C

sam (.sam.):

3/8 - 1/8 = 2/8 2/8 divided by 2 = 1/4